An airplane flew for 6 hours with a 22-km/h tail wind. the return flight against the same wind took 8 hours. find the speed of the plane in stil air. how do i set up this equation?
Answer by mananth(16946)

let the distance be y
speed of plane in still air be x
withwind speed = x+22
against wind speed = x-22
time with wind = 6 hours
time against wind = 8 hours
distance = speed & time
(x+22)6 =Y
(x-22)*8 y
(x+22)*6 = (x-22)*8
6x+132=8x-176
8x-6x=132+176
2x=308
/2
x= 154
Speed of plane in still air = 154 mph